find all solutions for x in the interval (0,2pi) for the equation sin x = -.9589
Answer by lwsshak3(11628)

find all solutions for x in the interval (0,2pi) for the equation sin x = -.9589
**
sin function<0 in quadrants III and IV
Using inverse function key on calculator:
sin x inverse(-.9589)=73.52º reference angle
x=180+73.52=253.52º
and x= 360-73.52=286.48º